OpenStudy (anonymous):

what is the simplified form of 12x/x^2+6x+9/4x^2/x^2-9

OpenStudy (anonymous):

A. 3(x+3)/x(x-3) B. 3(x-3)/x(x+3) C. 3x(x+3)/(x-3) D. (x-3)/3x(x+3)

OpenStudy (e.mccormick):

Hmmm... looks like they might be asking that.

OpenStudy (e.mccormick):

\[\frac{\frac{12x}{x^2+6x+9}}{\frac{4x^2}{x^2-9}}\]Short answer: Factor, do the division of the overall fraction, simplify. Which of those steps are you having problems with?